OpenStudy (anonymous):

The volume of a rectangular prism is represented by the function x3 + 9x2 + 6x – 16. The length of the box is x + 2 while the height is x + 8. Find the expression representing the width of the box.

OpenStudy (anonymous):

V = L w h\[w=\frac{V}{L h} \]\[\frac{x^3+9 x^2+6 x-16}{(x+2)(x+8) }=\frac{(x-1) (x+2) (x+8)}{(x+2)(x+8)}=x-1 \]